Night Plan UT Dec 01 2020


1st half PEPSi / 2nd half D/E time

Night Center 00:08 (local time)

Handover: 00:03 (split reconfiguration time)
  • Reset SX AdSec to 30 degree elevaton limit before proceeding
  • Reconfigure for LBCs - 20min
    • Filter focus offset tests - 2hrs - clear/thin skies - requires seeing 1" or better
    • TMS testing - 2hrs
    • non-lineariry/gain tests - 1 hr - requires photometric conditions
  • Configure for LUCI-Bino (authorize LUCI-1) - if unable to use LBCs
    • OSCO LUCI-1 (NCPA off and NCPA on) - 0.5 hr
    • Palomar-2 ESM observations (LUCI-1 ESM, LUCI-2 SL) - 1 hour - Note: 12-17 degree from Moon - if clear this should be OK (clear and 1" seeing or better conditions ideal for comparison with previous work)
      • AO Ref star is a double-star, but has been successfully closed on with FLAO
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/ESM_Testing
      • If any LUCI-2 issues, simply execute the LUCI-1 script in mono mode
      • If clear/photometric - observe P247 and G1919B2b before and after Pal2
    • If conditions are too poor to observe with LBC or proceed with Pal-2
      • M34 ESM (AO Ref R=12) - K,H,J filters - 0.75 hour - 30 deg from Moon
      • M67 ESM (AO Ref R=12, AO Ref=16.3) - K,H,J filters - 2 hour - > 64 deg from Moon
        • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/ESM_Testing
      • ICRF3 Poor weather backup - observe until targets exhausted
        • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/ICRF3_Rothberg_2020B
  • Link to A. Cardwell's night log on googledocs
Issue with the LBC swingarms - 2nd half of night switched to LUCI AO/ESM testing

-- BarryRothberg - 28 Nov 2020
Topic revision: r4 - 30 Nov 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback